OpenStudy (anonymous):

How many liters of a 40% acid solution must be added to 12 liters of a 20% solution r=to obtain a 25% solution?

OpenStudy (dumbcow):

set up a equation equating the amount of acid --> (.4)x + (.2)*12 = (.25)(x+12) solve for x

OpenStudy (dumbcow):

you can get rid of decimals by multiplying entire equation by 100 40x +20*12 = 25(x+12)